# The Quasar ULAS J1120+0641 May Never Reach a āNormalā Black Hole to Stellar Mass Ratio

### v1

@@ 5.2.2 Older, optically visible host galaxy
If the flux from the older stellar population (visible to JWST) is similarly obscured to the newly forming population, the mass will be underestimated; this may have contributed to the large positive uncertainty in the Prospector models. We can show more directly that this is unlikely to be a major issue by fitting the combined galaxy flux measurements from Stone etĀ al. (2024) and Marshall etĀ al. (2024) (reported in TableĀ 3). FigureĀ 6 shows these fits: models with a broad range of star formation histories can fit the data without invoking large extinction.
@@ 5.2.3 Potential total host galaxy mass
Combining the masses of the stars forming in the ULIRG process, the older stars, and in the merging companion discovered by Marshall etĀ al. (2024), the total potential stellar mass for the host galaxy becomes Mā, assuming a local, Chabrier/Kroupa IMF for the stellar population(s).
@@
Another contribution to the host galaxy stellar mass is its atomic and molecular gas reservoir, which does not contribute to its luminosity as observed by JWST. Values for the mass of this gas are vary significantly. Neeleman etĀ al. (2021) provide an estimate of Mā by converting the far-IR flux density from ALMA to a dust mass and assuming a dust-to-gas mass ratio. They also estimate a mass of Mā directly from the [CĀ II] luminosity. Venemans etĀ al. (2017) derive upper limits for the gas mass using two mm-wave lines with ALMA: from an upper limit on the CO(2-1) line flux; and Mā from an upper limit on the CO(7-6) line flux. All of these values are consistent with Mā, which we adopt. Only a fraction of this gas will eventually be converted into stars, governed by the star formation efficiency (SFE). Polzin etĀ al. (2024) find SFEs of a few %, independent of metallicity. Kim etĀ al. (2021) find possible SFEs of a few % to about 10%. Andalman etĀ al. (2024) run simulations for where the high gas density can favor increased SFE, and typically find values between 10% and 20%. However, estimates of the SFE are highly variable in general, depending on the assumed boundary conditions and approach. A reasonable number for J1120+0641 is 20%, in which case the gas presently in the galaxy could eventually form about Mā of stars.
@@
Star formation efficiency to | 0.10 | 0.20 | 0.40 | A24, K21, P24 | |
@@
Star formation efficiency to | 0.1 | 0.2 | 0.4 | A24, K21, P24 | |
@@
Host Stellar Mass at : | 1.53 | 5.99 | 25.8 | Mā |
@@ 5.3 The Future of ULAS J1120+0641
Under this choice, the narrow-band candidates across both fields have a total mass of approximately Mā ( Mā if the medium-confidence sources, two of which are fit to dubiously high masses, are included). Mā of galaxies lies in the quasar field. All galaxies in the quasar field are high-confidence sources (see Figure 2).
@@
With a robust estimate of the total stellar mass in the satellite galaxies, we finally need to account for the gas present in the satellite galaxies, some fraction of which will form stars. While no comprehensive surveys from e.g. ALMA exist at the high redshift and very low mass of our sample galaxies, galaxies of slightly higher mass at the same redshift tend to have gas masses approximately an order of magnitude larger than their stellar masses (Algera etĀ al., 2025; de Graaff etĀ al., 2024; Heintz etĀ al., 2022). Assuming a gas-to-stars ratio of 10 and a star formation efficiency of 20%, the total stellar mass of Mā in the surrounding galaxies implies that an additional Mā of stars can form by . Between the detected galaxies, the galaxies missed due to incompleteness, and their gas, we therefore conclude that approximately Mā of stars could possibly be added to the host galaxyās mass between its observed redshift and through mergers. This value is highly uncertain, dependent on the assumed gas-to-stars ratio and star formation efficiency: this mass from āfuture starsā dominates the uncertainty on our final host galaxy mass (see Table 4). To see the effect that different assumptions have on the derived total mass, see Table 4.
Combining the mass of the surrounding galaxies with that of the host gives a total mass of Mā. Assuming that no further black hole growth occurs, then the host of J1120+0641 will grow to , or 2.5%, at : still about a factor of 25 greater than the typical value for low-redshift galaxies of similar stellar mass (, Greene etĀ al., 2020). Only as we approach the upper end of our uncertainty rangeāstill assuming no further black hole growth, and adopting the large Prospector masses for the satellites as well as large gas fractions and very efficient star formationādo we achieve a final stellar mass of Mā, for (still on the lower end of the expected stellar mass given the mass of J1120+0641ās black hole).
@@ Appendix B Quasar luminosity
Our determination of the luminosity of J1120+0641 is centered on FigureĀ 11. The presentation shows relative luminosity for different wavelengths directly. The luminosity largely emerges between 0.05 and 30 m. The spectra and photometry cover a significant fraction of this wavelength range, except for m. The extinction to the BLR is minimal (Bosman etĀ al., 2024), so the SED short of 0.1 m should be reasonably accurate. We therefore estimate the luminosity for the normal and WDD cases Lyu etĀ al. (2017) by integrating the output for the appropriate SED, finding Lā and Lā, respectively for ānormalā and WDD. Extrapolating from Figure 12 in Lyu etĀ al. (2017), it appears that there is about a 30% probability that the WDD template is appropriate. That is, the most likely luminosity is Lā, or ergs s-1. .

### v3

@@ 5.2.2 Older, optically visible host galaxy
If the flux from the older stellar population (visible to JWST) is similarly obscured to the newly forming population, the mass will be underestimated; this may have contributed to the large positive uncertainty in the Prospector models. We can show more directly that this is unlikely to be a major issue by fitting the combined galaxy flux measurements from Stone etĀ al. (2024) and Marshall etĀ al. (2024) (reported in TableĀ 3). FigureĀ 6 shows these fits: models with a broad range of star formation histories can fit the data without invoking large extinction. In addition, there is no evidence from the photometry for extinction sufficiently strong to result in a significant underestimate of the galaxy mass.
@@ 5.2.3 Potential total host galaxy mass
Combining the Mā of stars forming in the ULIRG process (Section 5.2.1), the Mā of older stars (Section 5.2.2), and Mā in the merging companion discovered by Marshall etĀ al. (2024), the total potential stellar mass for the host galaxy becomes Mā, assuming a local, Chabrier/Kroupa IMF for the stellar population(s).
@@
Another contribution to the host galaxy stellar mass is its atomic and molecular gas reservoir, which does not contribute to its luminosity as observed by JWST. Values for the mass of this gas are vary significantly. Neeleman etĀ al. (2021) provide an estimate of Mā by converting the far-IR flux density from ALMA to a dust mass and assuming a dust-to-gas mass ratio. They also estimate a mass of Mā directly from the [CĀ II] luminosity. Venemans etĀ al. (2017) derive upper limits for the gas mass using two mm-wave lines with ALMA: from an upper limit on the CO(2-1) line flux; and Mā from an upper limit on the CO(7-6) line flux. All of these values are consistent with Mā, which we adopt. Only a fraction of this gas will eventually be converted into stars, governed by the star formation efficiency (SFE). Polzin etĀ al. (2024) find SFEs of a few %, independent of metallicity. Kim etĀ al. (2021) find possible SFEs of a few % to about 10%. Andalman etĀ al. (2025) run simulations for where the high gas density can favor increased SFE, and typically find values between 10% and 20%. However, estimates of the SFE are highly variable in general, depending on the assumed boundary conditions and approach. To be conservative and avoid underestimating the mass, we adopt a fiducial value of 20% for J1120+0641, in which case the gas presently in the galaxy could eventually form about Mā of stars. For the effect of other choices of SFE on the final mass, see Table 4.
@@
| Star formation efficiency to | 0.10 | 0.20 | 0.40 | An25, K21, P24 | |
@@
| Star formation efficiency to | 0.1 | 0.2 | 0.4 | An25, K21, P24 | |
@@
| Host Stellar Mass at : | 1.53 | 5.99 | 25.8 | Mā | |
@@ 5.3 The Future of ULAS J1120+0641
Under this choice, the narrow-band candidates across both fields have a total stellar mass of approximately Mā ( Mā if the medium-confidence sources, two of which are fit to dubiously high masses, are included). Mā of galaxies lies in the quasar field. All galaxies in the quasar field are high-confidence sources (see Figure 2).
@@
With a robust estimate of the total stellar mass in the satellite galaxies, we finally need to account for the gas present in the satellite galaxies, some fraction of which will form stars. While no comprehensive surveys from e.g. ALMA exist at the high redshift and very low mass of our sample galaxies, galaxies of slightly higher mass at the same redshift tend to have gas masses approximately an order of magnitude larger than their stellar masses (Algera etĀ al., 2025; de Graaff etĀ al., 2024; Heintz etĀ al., 2022). Assuming a gas-to-stars ratio of 10 and an SFE of 20% (see references in Section 5.2.3), the total stellar mass of Mā in the surrounding galaxies implies that an additional Mā of stars can form by . For the effect of changing the adopted gas-to-stars ratio and SFE on the final mass, see 4. Between the detected galaxies, the galaxies missed due to incompleteness, and their gas, we therefore conclude that approximately Mā of stars could possibly be added to the host galaxyās mass between its observed redshift and through mergers. This value is highly uncertain, dependent on the assumed gas-to-stars ratio and star formation efficiency: this mass from āfuture starsā dominates the uncertainty on our final host galaxy mass (see Table 4). To see the effect that different assumptions have on the derived total mass, see Table 4.
Combining the mass of the surrounding galaxies with that of the host gives a total mass of Mā. Assuming that no further black hole growth occurs, then the host of J1120+0641 will grow to , or 2.5%, at : still about a factor of 25 greater than the typical value for low-redshift galaxies of similar stellar mass (, Greene etĀ al., 2020). Only as we approach the upper end of our uncertainty rangeāstill assuming no further black hole growth, and adopting the large Prospector masses for the satellites as well as large gas fractions and very efficient star formationādo we achieve a final stellar mass of Mā, for (still on the lower end of the expected stellar mass given the mass of J1120+0641ās black hole). The range in that J1120+0641 could reach is shown in Figure 8.
@@ Appendix B Quasar luminosity
Our determination of the luminosity of J1120+0641 is centered on FigureĀ 12. The presentation shows relative luminosity for different wavelengths directly. The luminosity largely emerges between 0.05 and 30 m. The spectra and photometry cover a significant fraction of this wavelength range, except for m. The extinction to the BLR is minimal (Bosman etĀ al., 2024), so the SED short of 0.1 m should be reasonably accurate. We therefore estimate the luminosity for the normal and WDD cases Lyu etĀ al. (2017) by integrating the output for the appropriate SED, finding Lā and Lā, respectively for ānormalā and WDD. Extrapolating from Figure 12 in Lyu etĀ al. (2017), it appears that there is about a 30% probability that the WDD template is appropriate. That is, the most likely luminosity is Lā, or ergs s-1. .